OpenStudy (anonymous):

Choose the correct simplification of (6x^6y^12z^3)^0(2x)^3. Answer 6x^6y^12z^3 0 8x^3 48x^9y^12z^3

Parth (parthkohli):

Well actually anything to the power 0 is 1.

OpenStudy (anonymous):

so basically not even there

Parth (parthkohli):

Then the answer would be: \(\Large \color{Black}{\Rightarrow 1 \times (2x)^3 }\)

OpenStudy (anonymous):

?

Parth (parthkohli):

Because anything to the power 0 is 1, then (6x^6y^12z^3)^0 is simplified to 1.

OpenStudy (anonymous):

so the answer would be 0?

Parth (parthkohli):

No! I just told you that it's equal to 1 * (2x)^3

OpenStudy (anonymous):

because anything to the power 0 is 1? not really..

Parth (parthkohli):

Yes, anything to the power 0 is 1.

OpenStudy (anonymous):

I am confused :S

OpenStudy (anonymous):

\[0^0\] \[\infty^0\]

Parth (parthkohli):

Hmm yeah you're right about 0.

Parth (parthkohli):

But here most probably it's a non zero

OpenStudy (anonymous):

so whats the answer?

Parth (parthkohli):

1 * (2x)^3 = (2x)^3 = 8x^3

OpenStudy (anonymous):

k got it